# Rustam Ibragimova,b    Jihyun Kimc    Anton Skrobotovd a Imperial College Business School b New Economic School

### v2

@@ 5.2.2 Results
The results, summarized in Tables 611, indicate that the proposed tests exhibit excellent size control under the null hypothesis across all DGPs, whereas the IVX test is substantially oversized, particularly when volatility is nonstationary or exhibits structural breaks. Furthermore, both the hybrid and $t$-statistic approaches demonstrate nontrivial power, even though they are constructed using only half of the observations. Among the size-controlled procedures, the statistic $t_{16}(\tau(\check{\beta}_{o}))$ consistently delivers the strongest performance.
Overall, these findings corroborate the theoretical robustness of our methods. They remain valid under heavy-tailed, endogenous, and persistent regressors, as well as under heteroskedastic and persistent volatility. In contrast, the IVX test performs well only under stationary volatility and light tails. Hence, our proposed robust procedures offer a practical and reliable alternative to existing inference methods for predictive regressions in both continuous- and discrete-time frameworks.
@@ 6 Empirical Application
To illustrate the empirical performance of the proposed tests relative to the Cauchy RT and Cauchy VC tests, we reexamine the dataset used by CJP2016 to test the predictability of stock returns using the dividend–price (D/P) and earnings–price (E/P) ratios as predictors. For stock returns, we employ the NYSE/AMEX value-weighted index and the S&P 500 index obtained from the Center for Research in Security Prices (CRSP). The dividend–price ratio is defined as the annual dividend divided by the current total market value. Further details on data construction are provided in Section 6.1 of CJP2016.
Following CJP2016, we estimate two types of predictive regressions: one based on all returns and another based only on returns generated from the diffusive component of stock prices, obtained by first testing for jumps and removing observations corresponding to detected jumps. In all cases, we apply one-sided tests.
@@
Consistent with our simulation evidence, the Cauchy RT test demonstrates strong finite-sample power but requires high-frequency data due to its reliance on a continuous-time approximation. [^18] For the Cauchy RT test in our simulations, we estimate the discretized time-changed regression using $n$ lower-frequency observations, with $n=12T$ (approximately monthly), generated by the random time-sampling scheme described in Section 5 of CJP2016. The mixed empirical results—where the Cauchy RT test fails to reject the null while both the proposed methods and the Cauchy VC test do reject—may reflect the limited accuracy of the continuous-time approximation when applied to monthly, quarterly, or yearly data. Evaluating the robustness of the continuous-time approximation underlying the Cauchy RT test remains an interesting topic for future research.
@@ 7 Conclusion
Simulation evidence demonstrates that the proposed tests perform well in finite samples, maintaining correct size and competitive power under a wide range of data-generating processes, including those characterized by stochastic volatility, structural breaks, and regime switching. Although our procedures require the assumption of asymptotically exogenous volatility, they exhibit excellent robustness and complement existing Cauchy-based methods, including the IVX method of KMS2015, the Cauchy RT test of CJP2016 and the Cauchy VC test of ibragimov2023new.
@@
Part (a) follows from Theorem 1 and the discussion in Section 2.2 of IbragimovMuller2010. For part (b), we deduce from Lemma 3.1 that
$\frac{1}{\sqrt{T/q}}\check{\gamma}_{j}=\frac{1}{\sqrt{T/q}}\zeta_{j}+o_{p}\!\left(\frac{1}{\sqrt{T}}\right)\to_{p}\beta\,\mathbb{E}|x_{t}|$ (A.2)
@@
$\frac{\sqrt{q}\,\overline{\gamma}}{\sqrt{T}}\to_{p}\beta\,\mathbb{E}|x_{t}|.$ (A.3)
@@
$\mathbb{E}[z_{t-1}y_{t}]=\mathbb{E}\!\left[\operatorname{diag}(|x_{1,t-1}|,\ldots,|x_{K,t-1}|)\right]C.$
Given $0<\mathbb{E}[|x_{k,t-1}|]<\infty$ for all $k$, $H_{0}^{(k)}$ holds for all $k$ if and only if $\mathbb{E}[z_{t-1}y_{t}]=\mathbf{0}_{K\times 1}$.
@@
$\mathbb{E}[z_{t-1}y_{t}]=\mathbb{E}[z_{t-1}x_{t-1}^{\prime}]B,$
and since $\mathbb{E}[z_{t-1}x_{t-1}^{\prime}]$ is assumed invertible, the condition $\mathbb{E}[z_{t-1}y_{t}]=\mathbf{0}_{K\times 1}$ implies $B=0$, completing the proof. ∎
@@
By construction, $(\xi_{t})$ is an MDS with respect to $(\mathcal{F}_{t})$ and satisfies $\mathbb{E}(\xi_{t}^{2}\mid\mathcal{F}_{t-1})=v_{t}^{2}$ under Assumption 2.1. The desired convergence then follows directly from the martingale central limit theorem, given Assumptions 2.2 and 2.3. ∎
Table 1: Size for the continuous time models
@@
T 5 20 50 5 20 50 5 20 50
CNST Cauchy RT 5.3 4.9 5.3 5.2 5.4 4.7 5.5 5.1 5.1
Cauchy VC 5.6 5.0 5.3 5.4 5.0 5.1 5.4 5.0 4.8
$t_{8}(\check{\gamma})$ 4.8 4.6 5.2 4.7 5.0 5.1 4.8 4.9 5.0
$t_{12}(\check{\gamma})$ 4.7 4.7 5.0 4.7 5.1 4.9 5.0 5.2 4.9
$t_{16}(\check{\gamma})$ 4.8 4.8 5.1 4.9 5.0 4.9 4.7 4.9 4.9
$\tau({\check{\beta}})$ 5.2 4.8 5.2 5.0 4.8 5.1 5.2 4.9 4.8
SB Cauchy RT 5.6 5.0 5.1 5.2 5.3 5.0 5.4 5.0 4.9
Cauchy VC 8.0 6.7 6.3 7.8 6.5 6.0 7.9 6.4 6.0
$t_{8}(\check{\gamma})$ 3.6 3.7 3.9 4.0 4.1 3.7 3.8 3.7 3.7
$t_{12}(\check{\gamma})$ 4.7 4.2 4.6 4.3 4.6 4.5 4.2 4.2 4.5
$t_{16}(\check{\gamma})$ 4.7 4.6 4.6 4.8 4.7 4.5 4.3 4.6 4.5
$\tau({\check{\beta}})$ 6.1 5.0 4.9 6.1 5.5 5.1 6.2 5.5 5.0
RS Cauchy RT 5.0 4.8 5.2 4.9 4.9 4.9 5.4 5.1 4.8
Cauchy VC 5.2 5.4 6.1 5.2 5.1 5.8 5.6 5.8 5.8
$t_{8}(\check{\gamma})$ 4.5 4.5 5.1 4.6 4.4 4.9 4.6 5.3 4.6
$t_{12}(\check{\gamma})$ 4.8 4.6 5.0 4.5 4.8 4.8 4.6 4.9 4.7
$t_{16}(\check{\gamma})$ 5.0 4.4 4.9 4.6 4.6 4.9 5.1 5.1 4.5
$\tau({\check{\beta}})$ 5.0 4.9 5.2 5.2 4.7 4.9 5.2 5.3 5.0
GBM Cauchy RT 4.4 4.7 4.4 4.3 4.5 4.4 4.6 4.5 4.5
Cauchy VC 5.4 5.5 6.1 5.7 5.7 5.9 5.7 5.9 6.5
$t_{8}(\check{\gamma})$ 3.2 3.1 3.2 3.6 3.8 3.4 3.6 3.8 3.9
$t_{12}(\check{\gamma})$ 3.8 3.6 3.7 4.2 4.3 3.6 4.2 4.2 4.3
$t_{16}(\check{\gamma})$ 3.7 3.8 4.1 4.4 4.2 4.0 4.6 4.5 4.6
$\tau({\check{\beta}})$ 5.3 4.6 5.0 5.2 4.8 4.9 5.6 4.9 5.2
CNST, SB, GBM, and RS denote constant volatility, structural break, geometric Brownian motion, and regime switching, respectively.
Table 2: Power for the continuous time models, constant volatility
@@
Table 3: Power for the continuous time models, structural break in volatility
@@
Table 4: Power for the continuous time models, regime switching in volatility
@@
Table 5: Power for the continuous time models, geometric Brownian motion
@@
Table 6: Size and power for the discrete time models, constant volatility, MA(2) innovation
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 4.7 4.5 4.8 3.4 4.1 4.5 3.7 5.0 4.6
$t_{12}(\tau(\check{\beta}_{o}))$ 5.1 4.7 4.5 3.1 3.8 4.4 2.8 4.0 4.2
$t_{16}(\tau(\check{\beta}_{o}))$ 4.9 4.9 5.1 2.8 4.0 4.2 2.5 3.5 3.9
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 15.4 15.6 15.9 12.7 14.2 15.4 18.7 20.5 21.9
$t_{12}(\tau(\check{\beta}_{o}))$ 20.3 20.9 20.2 13.8 15.1 15.8 18.2 21.4 21.6
$t_{16}(\tau(\check{\beta}_{o}))$ 24.8 24.9 26.0 15.0 16.7 18.4 18.3 20.9 22.3
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 33.8 33.5 33.6 29.7 32.4 33.0 45.1 48.9 50.7
$t_{12}(\tau(\check{\beta}_{o}))$ 47.1 47.4 47.1 35.0 36.4 37.0 47.6 51.9 52.6
$t_{16}(\tau(\check{\beta}_{o}))$ 58.8 58.6 59.4 42.2 42.0 43.0 52.3 53.8 56.0
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 52.8 52.0 52.0 49.0 52.4 53.0 68.4 74.0 75.4
$t_{12}(\tau(\check{\beta}_{o}))$ 71.3 70.8 70.7 58.8 59.8 59.6 73.6 77.4 79.2
$t_{16}(\tau(\check{\beta}_{o}))$ 83.7 83.2 83.4 69.8 68.5 68.5 80.0 80.8 82.7
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 68.1 67.0 66.7 65.5 68.5 69.1 83.7 88.5 89.4
$t_{12}(\tau(\check{\beta}_{o}))$ 86.3 85.0 85.1 76.8 77.4 77.7 88.8 91.1 92.6
$t_{16}(\tau(\check{\beta}_{o}))$ 94.3 93.9 94.1 86.2 85.3 85.6 93.2 94.0 94.8
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 78.5 77.8 77.4 76.7 79.5 80.5 90.9 94.7 95.7
$t_{12}(\tau(\check{\beta}_{o}))$ 93.5 92.9 92.9 87.7 87.8 88.7 95.4 96.6 97.5
$t_{16}(\tau(\check{\beta}_{o}))$ 98.2 98.1 98.0 94.8 93.9 94.2 97.7 98.1 98.6
@@
Table 7: Size and power for the discrete time models, constant volatility, MA(4) innovation
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 5.0 4.9 4.8 5.2 5.3 4.8 4.5 4.7 5.2
$\tau(\check{\beta}_{e})$ 4.8 4.7 5.2 5.0 4.8 4.8 5.1 5.0 4.9
$t_{8}(\tau(\check{\beta}_{o}))$ 4.7 4.5 4.8 3.4 4.1 4.5 3.7 5.0 4.6
$t_{12}(\tau(\check{\beta}_{o}))$ 5.1 4.7 4.5 3.1 3.8 4.4 2.8 4.0 4.2
$t_{16}(\tau(\check{\beta}_{o}))$ 4.9 4.9 5.1 2.8 4.0 4.2 2.5 3.5 3.9
IVX 14.2 13.5 12.3 10.2 10.5 10.3 9.5 10.3 10.5
$\tau(\check{\beta}_{e})$ 7.0 7.0 7.2 12.4 11.5 12.3 16.8 17.2 17.5
$t_{8}(\tau(\check{\beta}_{o}))$ 11.5 11.2 12.0 9.1 10.6 11.4 13.0 14.1 15.0
$t_{12}(\tau(\check{\beta}_{o}))$ 14.4 14.4 14.1 9.6 10.9 11.7 11.6 14.1 14.5
$t_{16}(\tau(\check{\beta}_{o}))$ 16.7 16.5 17.6 10.1 11.7 13.0 11.3 13.6 14.6
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\tau(\check{\beta}_{e})$ 10.3 10.0 10.4 25.4 24.2 25.4 40.3 40.3 40.8
$t_{8}(\tau(\check{\beta}_{o}))$ 22.5 22.6 23.0 18.7 21.1 22.0 28.7 31.9 33.6
$t_{12}(\tau(\check{\beta}_{o}))$ 31.2 31.4 30.4 21.5 23.2 24.1 29.0 33.4 33.6
$t_{16}(\tau(\check{\beta}_{o}))$ 38.7 38.8 39.2 25.1 26.4 28.0 31.0 33.4 34.9
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\tau(\check{\beta}_{e})$ 14.4 14.4 14.4 42.8 42.5 42.8 65.9 66.9 66.5
$t_{8}(\tau(\check{\beta}_{o}))$ 36.4 36.0 36.0 32.2 34.7 35.2 48.3 52.5 54.5
$t_{12}(\tau(\check{\beta}_{o}))$ 50.3 50.5 50.3 37.8 39.1 40.0 51.2 55.2 56.5
$t_{16}(\tau(\check{\beta}_{o}))$ 62.5 62.3 63.3 46.0 45.4 46.1 56.9 57.9 60.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\tau(\check{\beta}_{e})$ 19.4 19.1 19.1 59.5 60.2 60.6 83.4 84.7 84.8
$t_{8}(\tau(\check{\beta}_{o}))$ 49.9 49.0 49.3 45.7 48.9 49.7 65.1 70.4 72.2
$t_{12}(\tau(\check{\beta}_{o}))$ 67.9 67.5 67.1 55.1 56.2 55.9 69.9 73.8 75.9
$t_{16}(\tau(\check{\beta}_{o}))$ 80.4 80.3 80.5 65.8 64.5 64.3 76.5 77.2 79.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\tau(\check{\beta}_{e})$ 24.5 24.0 24.6 73.7 74.3 75.5 92.8 94.1 94.1
$t_{8}(\tau(\check{\beta}_{o}))$ 61.6 60.4 60.7 58.5 61.6 62.0 77.8 83.2 84.4
$t_{12}(\tau(\check{\beta}_{o}))$ 80.4 79.2 79.4 69.3 70.5 70.1 83.1 86.1 87.8
$t_{16}(\tau(\check{\beta}_{o}))$ 90.6 90.1 90.3 79.9 78.8 79.0 88.7 89.3 90.8
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Table 8: Size and power for the discrete time models, structural break in volatility, MA(2) innovation
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 3.7 3.7 3.7 2.4 3.4 3.5 2.8 3.5 3.4
$t_{12}(\tau(\check{\beta}_{o}))$ 4.5 4.3 3.8 3.0 3.6 3.8 2.7 3.3 3.6
$t_{16}(\tau(\check{\beta}_{o}))$ 4.3 4.8 4.6 2.7 3.4 4.1 2.5 3.1 3.7
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 11.5 11.6 11.7 9.2 10.5 10.8 12.9 15.1 15.6
$t_{12}(\tau(\check{\beta}_{o}))$ 16.0 16.0 16.0 10.8 12.1 12.4 13.3 16.1 16.6
$t_{16}(\tau(\check{\beta}_{o}))$ 19.6 20.2 20.5 12.2 13.8 14.6 14.6 16.1 17.6
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 24.8 24.7 24.8 21.1 23.1 23.4 32.0 35.3 36.3
$t_{12}(\tau(\check{\beta}_{o}))$ 36.3 36.6 36.4 27.3 28.2 28.0 35.0 39.3 39.6
$t_{16}(\tau(\check{\beta}_{o}))$ 44.8 45.7 46.1 31.5 33.6 34.2 39.5 41.0 43.0
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 39.4 38.6 39.3 35.6 37.6 38.0 51.1 56.1 57.6
$t_{12}(\tau(\check{\beta}_{o}))$ 56.3 57.2 56.4 45.7 46.4 45.9 57.6 62.3 63.3
$t_{16}(\tau(\check{\beta}_{o}))$ 67.5 69.2 69.3 53.9 55.3 55.1 63.8 65.9 67.7
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 51.7 51.4 51.6 48.5 50.0 51.8 65.0 71.3 72.7
$t_{12}(\tau(\check{\beta}_{o}))$ 71.7 72.5 71.4 62.0 62.2 62.1 74.0 77.3 79.1
$t_{16}(\tau(\check{\beta}_{o}))$ 82.3 83.1 83.1 71.6 72.3 72.0 81.1 82.0 83.9
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 61.7 61.1 61.2 58.6 60.4 62.5 74.1 80.5 82.4
$t_{12}(\tau(\check{\beta}_{o}))$ 81.7 82.3 81.3 73.6 73.7 74.4 84.0 86.5 88.2
$t_{16}(\tau(\check{\beta}_{o}))$ 90.6 91.1 91.4 83.2 83.1 83.0 90.1 91.1 91.9

### v3

@@ 5.2.2 Results
The results, summarized in Tables S.98, indicate that the proposed tests exhibit excellent size control under the null hypothesis across all DGPs, whereas the IVX test is substantially oversized, particularly when volatility is nonstationary or exhibits structural breaks. Furthermore, the hybrid approaches demonstrate nontrivial power, even though they are constructed using only half of the observations.
Overall, these findings corroborate the theoretical robustness of our methods. They remain valid under heavy-tailed, endogenous, and persistent regressors, as well as under heteroskedastic and persistent volatility. In contrast, the IVX test exhibits severe size distortions in one-sided tests across all cases, including the constant volatility setting, as documented by Demetrescu et al. (2023).
We also conduct two-sided tests and report the results in the Online Appendix (Section B.2). In this case, the IVX method exhibits good size performance under constant volatility but suffers from severe size distortions in settings with time-varying volatility, whereas our methods maintain accurate size control across all specifications.
Taken together, these results suggest that the proposed robust procedures provide a practical and reliable alternative to existing inference methods for predictive regressions in both continuous- and discrete-time frameworks.
@@ 6 Empirical Application
To illustrate the empirical performance of the proposed tests relative to the Cauchy RT and Cauchy VC tests, we reexamine the dataset used by Choi et al. (2016) to test the predictability of stock returns using the dividend–price (D/P) and earnings–price (E/P) ratios as predictors. [^22] Further empirical results, including intercept-robust estimations and Wald-type tests compared to IVX (Kostakis et al., 2015), are detailed in the Online Appendix C. For stock returns, we employ the NYSE/AMEX value-weighted index and the S&P 500 index obtained from the Center for Research in Security Prices (CRSP). The dividend–price ratio is defined as the annual dividend divided by the current total market value. Further details on data construction are provided in Section 6.1 of Choi et al. (2016).
Following Choi et al. (2016), we estimate two types of predictive regressions: one based on all returns and another based only on returns generated from the diffusive component of stock prices, obtained by first testing for jumps and removing observations corresponding to detected jumps. In all cases, we apply one-sided tests.
@@
Consistent with our simulation evidence, the Cauchy RT test demonstrates strong finite-sample power but requires high-frequency data due to its reliance on a continuous-time approximation. [^23] For the Cauchy RT test in our simulations, we estimate the discretized time-changed regression using $n$ lower-frequency observations, with $n=12T$ (approximately monthly), generated by the random time-sampling scheme described in Section 5 of Choi et al. (2016). The mixed empirical results—where the Cauchy RT test fails to reject the null while both the proposed methods and the Cauchy VC test do reject—may reflect the limited accuracy of the continuous-time approximation when applied to monthly, quarterly, or yearly data. Evaluating the robustness of the continuous-time approximation underlying the Cauchy RT test remains an interesting topic for future research.
@@ 7 Conclusion
Simulation evidence demonstrates that the proposed tests perform well in finite samples, maintaining correct size and competitive power under a wide range of data-generating processes, including those characterized by stochastic volatility, structural breaks, and regime switching. Although our procedures require the assumption of asymptotically exogenous volatility, they exhibit excellent robustness and complement existing methods, including the IVX method of Kostakis et al. (2015), the Cauchy RT test of Choi et al. (2016), and the Cauchy VC test of Ibragimov et al. (2024).
@@
Part (a) follows from Theorem 1 and the discussion in Section 2.2 of Ibragimov and Müller (2010). For part (b), we deduce from Lemma 3.1 that
$\frac{1}{\sqrt{T/q}}\check{\gamma}_{j}=\frac{1}{\sqrt{T/q}}\zeta_{j}+o_{p}\!\left(\frac{1}{\sqrt{T}}\right)\to_{p}\beta\,E|x_{t}|$ (A.2)
@@
$\frac{\sqrt{q}\,\overline{\gamma}}{\sqrt{T}}\to_{p}\beta\,E|x_{t}|.$ (A.3)
@@
$E[z_{t-1}y_{t}]=E\!\left[\operatorname{diag}(|x_{1,t-1}|,\ldots,|x_{K,t-1}|)\right]C.$
Given $0<E[|x_{k,t-1}|]<\infty$ for all $k$, $H_{0}^{(k)}$ holds for all $k$ if and only if $E[z_{t-1}y_{t}]=\mathbf{0}_{K\times 1}$.
@@
$E[z_{t-1}y_{t}]=E[z_{t-1}x_{t-1}^{\prime}]B,$
and since $E[z_{t-1}x_{t-1}^{\prime}]$ is assumed invertible, the condition $E[z_{t-1}y_{t}]=\mathbf{0}_{K\times 1}$ implies $B=0$, completing the proof. ∎
@@
By construction, $(\xi_{t})$ is an MDS with respect to $(\mathcal{F}_{t})$ and satisfies $E(\xi_{t}^{2}\mid\mathcal{F}_{t-1})=v_{t}^{2}$ under Assumption 2.1. The desired convergence then follows directly from the martingale central limit theorem, given Assumptions 2.2 and 2.3. ∎
Table 1: Size for Continuous-Time Models
@@
T 20 50 100 20 50 100 20 50 100
CNST Cauchy RT 4.9 5.3 5.0 5.4 4.7 5.0 5.1 5.1 5.1
Cauchy VC 5.0 5.3 4.9 5.0 5.1 5.0 5.0 4.8 5.0
$t_{8}(\check{\gamma})$ 4.6 5.2 4.9 5.0 5.1 5.4 4.9 5.0 5.2
$t_{12}(\check{\gamma})$ 4.7 5.0 5.1 5.1 4.9 5.4 5.2 4.9 5.0
$t_{16}(\check{\gamma})$ 4.8 5.1 5.0 5.0 4.9 5.3 4.9 4.9 5.1
$\tau(\check{\beta})$ 4.5 5.2 5.0 4.8 5.1 5.2 4.9 4.8 5.2
SB Cauchy RT 5.0 5.1 5.8 5.3 5.0 5.9 5.0 4.9 5.9
Cauchy VC 6.7 6.3 4.7 6.5 6.0 4.9 6.4 6.0 5.0
$t_{8}(\check{\gamma})$ 3.7 3.9 3.6 4.1 3.7 4.1 3.7 3.7 3.9
$t_{12}(\check{\gamma})$ 4.2 4.6 4.2 4.6 4.5 4.3 4.2 4.5 4.1
$t_{16}(\check{\gamma})$ 4.6 4.6 4.6 4.7 4.5 4.7 4.6 4.5 4.8
$\tau(\check{\beta})$ 5.0 4.9 4.9 5.5 5.1 5.0 5.5 5.0 4.8
RS Cauchy RT 4.8 5.2 6.6 4.9 4.9 6.5 5.1 4.8 6.2
Cauchy VC 5.4 6.1 5.1 5.1 5.8 5.2 5.8 5.8 4.7
$t_{8}(\check{\gamma})$ 4.5 5.1 4.6 4.4 4.9 5.1 5.3 4.6 4.5
$t_{12}(\check{\gamma})$ 4.6 5.0 4.7 4.8 4.8 5.3 4.9 4.7 5.1
$t_{16}(\check{\gamma})$ 4.4 4.9 4.7 4.6 4.9 5.2 5.1 4.5 5.0
$\tau(\check{\beta})$ 4.9 5.2 4.9 4.7 4.9 5.1 5.3 5.0 5.3
GBM Cauchy RT 4.7 4.4 6.6 4.5 4.4 6.5 4.5 4.5 6.4
Cauchy VC 5.5 6.1 4.4 5.7 5.9 4.8 5.9 6.5 5.0
$t_{8}(\check{\gamma})$ 3.1 3.2 3.2 3.8 3.4 3.5 3.6 3.9 3.9
$t_{12}(\check{\gamma})$ 3.3 3.7 3.8 4.3 3.6 4.0 4.2 4.3 4.4
$t_{16}(\check{\gamma})$ 3.8 4.1 4.1 4.2 4.0 4.2 4.5 4.6 4.3
$\tau(\check{\beta})$ 4.6 5.0 4.5 4.8 4.9 4.8 4.9 5.2 5.1
This table reports the rejection rates of the Cauchy RT, Cauchy VC, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and $\tau(\check{\beta})$ tests based on simulated data under $\beta=0$ (two-sided tests). The data are generated from the continuous-time model (11) with persistence parameter $\bar{\kappa}\in\{0,5,20\}$. The model is sampled at $\Delta=1/252$ (daily), so that a sample of length $T$ years contains $252T$ observations, with $T\in\{20,50,100\}$. CNST, SB, GBM, and RS denote constant volatility, structural breaks, geometric Brownian motion, and regime switching, respectively.
Table 2: Power for Continuous-Time Models with Constant Volatility
@@
This table reports the rejection rates of the Cauchy RT, Cauchy VC, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and $\tau(\check{\beta})$ tests based on simulated data under $\beta\in\{0.004,0.008,0.012,0,016,0.02\}$ (two-sided tests). The data are generated from the continuous-time model (11) with persistence parameter $\bar{\kappa}\in\{0,5,20\}$. The model is sampled at $\Delta=1/252$ (daily), so that a sample of length $T$ years contains $252T$ observations, with $T\in\{20,50,100\}$. The volatility is constant.
Table 3: Power for Continuous-Time Models with Structural Breaks in Volatility
@@
This table reports the rejection rates of the Cauchy RT, Cauchy VC, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and $\tau(\check{\beta})$ tests based on simulated data under $\beta\in\{0.004,0.008,0.012,0,016,0.02\}$ (two-sided tests). The data are generated from the continuous-time model (11) with persistence parameter $\bar{\kappa}\in\{0,5,20\}$. The model is sampled at $\Delta=1/252$ (daily), so that a sample of length $T$ years contains $252T$ observations, with $T\in\{20,50,100\}$. The volatility exhibits structural breaks.
Table 4: Power for Continuous-Time Models with Regime Switching in Volatility
@@
This table reports the rejection rates of the Cauchy RT, Cauchy VC, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and $\tau(\check{\beta})$ tests based on simulated data under $\beta\in\{0.004,0.008,0.012,0,016,0.02\}$ (two-sided tests). The data are generated from the continuous-time model (11) with persistence parameter $\bar{\kappa}\in\{0,5,20\}$. The model is sampled at $\Delta=1/252$ (daily), so that a sample of length $T$ years contains $252T$ observations, with $T\in\{20,50,100\}$. The volatility exhibits regime switching.
Table 5: Power for Continuous-Time Models with a Geometric Brownian Motion
@@
This table reports the rejection rates of the Cauchy RT, Cauchy VC, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and $\tau(\check{\beta})$ tests based on simulated data under $\beta\in\{0.004,0.008,0.012,0,016,0.02\}$ (two-sided tests). The data are generated from the continuous-time model (11) with persistence parameter $\bar{\kappa}\in\{0,5,20\}$. The model is sampled at $\Delta=1/252$ (daily), so that a sample of length $T$ years contains $252T$ observations, with $T\in\{20,50,100\}$. The volatility follows a geometric Brownian motion.
Table 6: Size and Power for Discrete-Time Models with Constant Volatility and MA(1) Innovations
@@
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (one-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(1) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility is constant.
Table 7: Size and Power for Discrete-Time Models with Structural Breaks in Volatility and MA(1) Innovations